the sides of a triangle have lengths 7, 8 and 10. what is the length of the shortest side of a similar triangle whose perimeter is 75
Answer by jorel1380(3719)

7+8+10=25
75/25=3
3(7)+3(8)+3(10)=21+24+30=75
The shortest side is 21..